OpenStudy (anonymous):

Q-12 Find the angle of elevation of the sun when a 6m high pole makes shadow of length 23m on the horizontal plane. a. 300 b. 450 c. 600 d. 900

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2.3 ?

OpenStudy (anonymous):

you can do your own homework now lols

OpenStudy (anonymous):

you just asked the same type of question

OpenStudy (radar):

I believe see was trying to express:\[2\sqrt{3}\] in this post. I base this because it was asked on previous post. I also believe see has questions or some confusion regarding the problem. First and foremost see must understand what is given! See must focus on the problem both what is available from information and secondly what he can reason from the problem.

OpenStudy (radar):

The diagram provided by elecengineer is something see needs to see (visualize). See you must visualize the sun and the sunbeam reaching down across that 6 meter pole creating a shadow. A shadow that is \[2\sqrt{3} \]meters in length. Also and very importantly see the triangle formed by the pole, earth surface with shadow, and of course see the imaginary ray grazing the top of the pole and terminating at the end of the shadow 2 sq rt 3 meters in length from the base of the pole.

OpenStudy (radar):

Maybe this also is a problem for see, the angle formed by this ray grazing the pole and terminating at the shadow's end point. This is the angle for which the problem is wanting solved. The tangent of that angle is the opposite side (pole 6 meter) divided by the length of the shadow of 2 sqrt3:\[\tan \theta=6/2\sqrt{3}\] Some additional math is needed to process this:\[6\over 2\sqrt{3}\]\[3\over \sqrt{3}\]\[3\times \sqrt{3}\over \sqrt{3}\times \sqrt{3}\]\[3\times \sqrt{3}\over3\]\[\sqrt{3}\] \[\tan^{-1} \sqrt{3}=60^{0}\]

OpenStudy (radar):

Presumably choice c. However, just selecting choice c and moving on, will not mean a successful career in a technical field, where a good understanding of math is a necessity. Study see, and learn.

OpenStudy (radar):

Or maybe study, learn and see!